# Estimating spectral radius of Bell-type operator via finite dimensional approximation of orthogonal projections

### v1

@@ 5.2 Equality conditions for the spectral radius estimates
First, we prove some properties of the eigenvalues of $A_{n}(\theta)+B_{n}(\theta)$ for $n\in\mathbb{N}$.
Lemma 5.4.
@@
    $\#(\sigma(A_{n}(\theta)+B_{n}(\theta))\cap 2\cos(I_{2n,k})\sin\theta)\leq 1.$
@@
We will first prove (vi) and (vii). In order to do so, we apply Yuen’s result [16], which is introduced in Section 2.4.2, to the symmetric tridiagonal matrix
@@
Let $a,b,c,\alpha,\beta,\gamma,\delta,\mu$ be complex numbers defined in Section 2.4.2. In the present setting,
@@
Then, for any $n\in\mathbb{N}$ and any $\lambda\in A_{n}(\theta)+B_{n}(\theta)$, there exists $\phi\in[0,2\pi)$ such that
$\lambda=\left\{\begin{array}[]{ll}2\sin\theta\cos\phi&\quad\sin\phi\neq 0,\\ 2\sin\theta&\quad\sin\phi=0,\end{array}\right$.
and $\phi$ satisfies
$\sin^{2}\theta\sin((2n+1)\phi)-(1+\cos\theta)^{2}\sin(2n-1)\phi=0.$
From Fact 2.5, the eigenvalues of $A_{n}(\theta)+B_{n}(\theta)$ are distinct for any $n\in\mathbb{N}$. Moreover, from statements (iv) and (v), statements (vi) and (vii) hold.
@@
$=-\dfrac{2u_{1}i}{\sqrt{\omega}}(1+\cos\theta)(\sin((j-1)\phi)-r\sin(j\phi))\quad(j=1,2,\dots,2n).$ (5.1)
@@
Example 5.5.
The following statements hold:
  1. (i)
    The spectrum of $A(\theta)+B(\theta)$ is given by
    $(-2\sin\theta,2\sin\theta)\subset F(A(\theta)+B(\theta))\subset\sigma(A(\theta)+B(\theta))=[-2\sin\theta,2\sin\theta].$
  2. (ii)
    The following equality holds:
    where $\lambda_{n}$ is the eigenvalue of $A_{n}(\theta)+B_{n}(\theta)$ whose absolute value is closest to $\sqrt{2}$ for $n\in\mathbb{N}$.
  3. (iii)
  4. (iv)
    We have
Proof of (i).
Let $\lambda\in(-2\sin\theta,2\sin\theta)$. From statements (vi) and (vii) of Lemma 5.4, there exists $\{\eta_{n}\}_{n=1}^{\infty}\subset(-2\sin\theta,2\sin\theta)$ and $\{u_{n}\}_{n=1}^{\infty}$ such that $\lim_{n\to\infty}\eta_{n}=\lambda$, and for any $n\in\mathbb{N}$, $u_{n}$ is a unit eigenvector of $A_{n}(\theta)+B_{n}(\theta)$ corresponding to $\eta_{n}$. Then for any $n\in\mathbb{N}$, $P_{n}(\theta)u_{n}=P(\theta)u_{n},\|(Q_{n}(\theta)-Q(\theta))u_{n}\|\leq 2|u_{n,2n}|$, where $u_{n,2n}$ is the $2n$-th component of $u_{n}$. From statement (viii) of Lemma 5.4, we have $\lim_{n\to\infty}\|(Q_{n}(\theta)-Q(\theta))u_{n}\|=0$. It follows that $\lambda\in F(A(\theta)+B(\theta))$. Hence we have $(-2\sin\theta,2\sin\theta)\subset F(A(\theta)+B(\theta))$. Since $\sigma(A(\theta)+B(\theta))$ is closed and $F(A(\theta)+B(\theta))\subset\sigma(A(\theta)+B(\theta))$, we conclude $2\sin\theta,2\sin\theta]\subset\sigma(A(\theta)+B(\theta))$.
We will prove that $\sigma(A(\theta)+B(\theta))\subset[-2\sin\theta,2\sin\theta]$. Let $\lambda\in\sigma(A(\theta)+B(\theta))$. Since $A(\theta)+B(\theta)$ is self-adjoint, $\lambda$ is in the approximate point spectrum of $A(\theta)+B(\theta)$. Therefore, there exists a sequence $\{u_{i}\}_{i=1}^{\infty}$ of unit vectors in $V$ such that $\|(A(\theta)+B(\theta))u_{i}-\lambda u_{i}\|<\dfrac{1}{2i}$ for any $i\in\mathbb{N}$. Since $\{(V_{i},\{P_{i}(\theta),Q_{i}(\theta)\})\}_{i=1}^{\infty}$ is a finite dimensional approximation of $(\ell^{2},\{P(\theta),Q(\theta)\})$, there exists a sequence $\{v_{i}\}_{i=1}^{\infty}$ of unit vectors in $V$ and exists a sequence $\varphi(1)<\ldots<\varphi(i)<\ldots$ of natural numbers such that $v_{i}\in V_{\varphi(i)}$ and $\|(A(\theta)+B(\theta))v_{i}-\lambda v_{i}\|<\dfrac{1}{i}$ for any $i\in\mathbb{N}$. It follows that $((A(\theta)+B(\theta))v_{i},v_{i})-\dfrac{1}{i}\leq\lambda\leq((A(\theta)+B(\theta))v_{i},v_{i})+\dfrac{1}{i}$ for any $i\in\mathbb{N}$. From Lemma 5.4, for any $i\in\mathbb{N}$
$((A(\theta)+B(\theta))v_{i},v_{i})=\lim_{n\to\infty}((A_{n}(\theta)+B_{n}(\theta))v_{i},v_{i})\in[-2\sin\theta,2\sin\theta].$
Thus, $2\sin\theta-\dfrac{1}{i}\leq\lambda\leq 2\sin\theta+\dfrac{1}{i}$ for any $i\in\mathbb{N}$. Letting $i\to\infty$, we have $2\sin\theta\leq\lambda\leq 2\sin\theta$. ∎
Proof of (ii).
From statements (vi)-(viii) of Lemma 5.4, we have
@@
Proof of (iii).
This follows directly from statement (i) and the fact that for any unitary representation $(\tau,V)$ of $D_{\infty}$ the spectrum of $\tau(g_{1})+\tau(g_{2})$ is invariant under any unitary intertwining operator. ∎
Proof of (iv).
@@
5.3 Application: Estimating spectral radius of Bell-CHSH operator
For bipartite quantum systems, we use the following notation.
Notation 5.6.
Let
  1. (S1)
    $V_{1}$ and $V_{2}$ be separable Hilbert spaces,
  2. (S2)
    $P_{i,j}$ be orthogonal projections on $V_{i}$ for $i,j\in\{1,2\}$,
  3. (A1)
  4. (S3)
  5. (S4)
    for $i\in\{1,2\}$, when $(V_{i},\{P_{i,1},P_{i,2}\})$ satisfies Case 1, let $\mathcal{A}_{i}$ be the approximation defined in Lemma 5.2,
  6. (S5)
  1. (i)
  2. (ii)
    $A_{i,n}=2P_{1,i,n}-\operatorname{id}_{V_{1,n}}$,
  3. (iii)
    $B_{i,n}=2P_{2,i,n}-\operatorname{id}_{V_{2,n}}$,
  4. (iv)
    $\lambda_{1,n}$ is the eigenvalue of $A_{1,n}+A_{2,n}$ whose absolute value is closest to $\sqrt{2}$,
  5. (v)
    $\lambda_{2,n}$ is the eigenvalue of $B_{1,n}+B_{2,n}$ whose absolute value is closest to $\sqrt{2}$.
Theorem 5.7.
Assume the setups (S1)-(S5), assumption (A1), notation (i)-(v) in Notation 5.6. Under these assumptions, we have
Acknowledgements
Y. F. would like to acknowledge Koichi Tojo, Takayuki Okuda, Yoji Koyama, Kohei Hatano, and Yuichiro Tanaka for their valuable comments and suggestions.

### v2

@@ 5.2 Equality conditions for the spectral radius estimates
Example 5.4.
The following statements hold:
  1. (i)
    The set $F(A(\theta)+B(\theta))$ satisfies
    $(0,2\sin\theta)\subset F(A(\theta)+B(\theta))\subset(0,2\sin\theta].$
  2. (ii)
    The following equality holds:
    where $\lambda_{n}$ is an eigenvalue of $A_{n}(\theta)+B_{n}(\theta)$ whose square is closest to $2$ for $n\in\mathbb{N}$.
  3. (iii)
  4. (iv)
    We have
The proof of this example is given in A.
5.3 Application: Estimating spectral radius of Bell-CHSH operator
For bipartite quantum systems, we use the following notation.
Notation 5.5.
Let
  1. (S1)
    $V_{1}$ and $V_{2}$ be separable Hilbert spaces,
  2. (S2)
    $P_{i,j}$ be orthogonal projections on $V_{i}$ for $i,j\in\{1,2\}$,
  3. (A1)
  4. (S3)
  5. (S4)
    for $i\in\{1,2\}$, when $(V_{i},\{P_{i,1},P_{i,2}\})$ satisfies Case 1, let $\mathcal{A}_{i}$ be the approximation defined in Notation 5.2,
  6. (S5)
  1. (i)
  2. (ii)
    $A_{i,n}=2P_{1,i,n}-\operatorname{id}_{V_{1,n}}$,
  3. (iii)
    $B_{i,n}=2P_{2,i,n}-\operatorname{id}_{V_{2,n}}$,
  4. (iv)
    $\lambda_{1,n}$ is an eigenvalue of $A_{1,n}+A_{2,n}$ whose square is closest to $2$,
  5. (v)
    $\lambda_{2,n}$ is an eigenvalue of $B_{1,n}+B_{2,n}$ whose square is closest to $2$.
Theorem 5.6.
Assume the setups (S1)-(S5), assumption (A1), notation (i)-(v) in Notation 5.5. Under these assumptions, we have
Acknowledgements
We thank Professors Albrecht Böttcher and Ilya M. Spitkovsky for their helpful comments and for sharing valuable insights relevant to this work. They kindly pointed out minor errors in the earlier version of our preprint [14] and helped us correct them, for which we are deeply grateful. Their observations greatly improved both the accuracy and the clarity of the current manuscript.
We also thank the anonymous referee for the careful reading of our manuscript and for the insightful comments that significantly improved the exposition and helped us better emphasize the contribution of our decomposition formula.
Y. F. would like to acknowledge Koichi Tojo, Takayuki Okuda, Yuichiro Tanaka, Yoji Koyama, and Kohei Hatano for their valuable comments and suggestions.
Appendix A Proof of Example 5.4
As a preparation, we prove some properties of the eigenvalues of $A_{n}(\theta)+B_{n}(\theta)$ for $n\in\mathbb{N}$.
Lemma A.1.
@@ Acknowledgements
    $\#(\sigma(A_{n}(\theta)+B_{n}(\theta))\cap 2\cos(I_{2n,k})\sin\theta)=1.$
@@
  9. (ix)
    Assume that $\theta<\dfrac{1}{2}\pi$. Let $\varepsilon>0$. Then for any sufficiently large $n\in\mathbb{N}$, there exists $\lambda\in[2-\varepsilon,2]$ such that
    $\sigma(A_{n}(\theta)+B_{n}(\theta))\cap[-2\sin\theta,2\sin\theta]^{c}=\{-\lambda,\lambda\}.$
  10. (x)
    Assume that $\theta>\dfrac{1}{2}\pi$. Then for any sufficiently large $n\in\mathbb{N}$,
    $\#(\sigma(A_{n}(\theta)+B_{n}(\theta))\cap 2\cos(I_{2n,0})\sin\theta)=1,$
    $\#(\sigma(A_{n}(\theta)+B_{n}(\theta))\cap 2\cos(I_{2n,2n})\sin\theta)=1,$
    $\sigma(A_{n}(\theta)+B_{n}(\theta))\subset[-2\sin\theta,2\sin\theta].$
@@
We will first prove (vi) and (vii). In order to do so, we apply the result of Yueh and Cheng [18], which is introduced in Section 2.4.2, to the symmetric tridiagonal matrix
@@
Let $a,b,c,\alpha,\beta,\gamma,\delta,\mu$ be complex numbers defined in Section 2.4.2. Let $n\in\mathbb{N}$. In the present setting,
@@
Then, if we find $\phi\in[0,2\pi)$ such that $\sin\phi\neq 0$ and $\phi$ satisfies
$\sin^{2}\theta\sin((2n+1)\phi)-(1+\cos\theta)^{2}\sin(2n-1)\phi=0,$
$\lambda=2\sin\theta\cos\phi\in\sigma(A_{n}(\theta)+B_{n}(\theta))$.
From Proposition 2.5, the eigenvalues of $A_{n}(\theta)+B_{n}(\theta)$ are distinct for any $n\in\mathbb{N}$. Moreover, from statements (iv) and (v), statements (vi) and (vii) hold.
@@
$=-\dfrac{2u_{1}i}{\sqrt{\omega}}(1+\cos\theta)(\sin((j-1)\phi)-r\sin(j\phi))\quad(j=1,2,\dots,2n).$ (A.1)
@@
Proof of (ix).
In Theorem 4 in [15], it is shown that $2\in\sigma(A(\theta)+B(\theta))$. Hence, there exists a unit vector $u\in\ell^{2}$ such that $|((A(\theta)+B(\theta))u,u)|>2-\dfrac{\varepsilon}{2}$. From Proposition 4.2, for any sufficiently large $n\in\mathbb{N}$ there exists a unit vector $u_{n}\in\mathbb{C}^{2n}$ such that $|((A_{n}(\theta)+B_{n}(\theta))u_{n},u_{n})|>2-\varepsilon$. Fix such $n\in\mathbb{N}$. Then $(A_{n}(\theta)+B_{n}(\theta))$ has an eigenvalue $\lambda\in[-2,-2+\varepsilon]\cup[2-\varepsilon,2]$.
From (vi), $(\sigma(A_{n}(\theta)+B_{n}(\theta))\cap[-2\sin\theta,2\sin\theta])\geq 2n-2$. Moreover, from Proposition 2.2 and Proposition 2.3, $\gamma\in(\sigma(A_{n}(\theta)+B_{n}(\theta))\cap[-2\sin\theta,2\sin\theta])$ for any $\gamma\in(\sigma(A_{n}(\theta)+B_{n}(\theta))\cap[-2\sin\theta,2\sin\theta])$. Since $\operatorname{tr}(A_{n}(\theta)+B_{n}(\theta))=0$, $\lambda\in A_{n}(\theta)+B_{n}(\theta)$. Consequently, we have
$\sigma(A_{n}(\theta)+B_{n}(\theta))\cap[-2\sin\theta,2\sin\theta]^{c}=\{-\lambda,\lambda\}.$
Proof of (x).
Note that
$\lim_{n\to\infty}f_{2n}\left(\dfrac{\pi}{2(2n+1)}\right)=\lim_{n\to\infty}f_{2n}\left(\pi-\dfrac{\pi}{2(2n+1)}\right)=1,$
and for any $n\in\mathbb{N}$
$\lim_{\phi\to\frac{\pi}{2n+1}-0}f_{2n}(\phi)=\lim_{\phi\to\left(\pi-\frac{\pi}{2n+1}\right)+0}f_{2n}(\phi)=\infty.$
Since for any $\theta\in\left(\dfrac{\pi}{2},\pi\right)$
$\dfrac{\sin^{2}\theta}{(1+\cos\theta)^{2}}=\tan^{2}\left(\dfrac{\theta}{2}\right)>1,$
from (iii), we have that for any sufficiently large $n\in\mathbb{N}$
$1\leq\min\{\#(Z_{2n}\cap I_{2n,0}),\#(Z_{2n}\cap I_{2n,2n})\}.$
Fix such $n\in\mathbb{N}$. Combining this and (vi), it follows that
$\sigma(A_{n}(\theta)+B_{n}(\theta))\subset[-2\sin\theta,2\sin\theta].$
We give the proof of Example 5.4 as follows.
Proof of (i) of Example 5.4.
Let $\lambda\in(0,2\sin\theta)$. From statements (vi) and (vii) of Lemma A.1, there exists $\{\eta_{n}\}_{n=1}^{\infty}\subset(-2\sin\theta,2\sin\theta)$ and $\{u_{n}\}_{n=1}^{\infty}$ such that $\lim_{n\to\infty}\eta_{n}=\lambda$, and for any $n\in\mathbb{N}$, $u_{n}$ is a unit eigenvector of $A_{n}(\theta)+B_{n}(\theta)$ corresponding to $\eta_{n}$. Then for any $n\in\mathbb{N}$, $P_{n}(\theta)u_{n}=P(\theta)u_{n},\|(Q_{n}(\theta)-Q(\theta))u_{n}\|\leq 2|u_{n,2n}|$, where $u_{n,2n}$ is the $2n$-th component of $u_{n}$. From statement (viii) of Lemma A.1, we have $\lim_{n\to\infty}\|(Q_{n}(\theta)-Q(\theta))u_{n}\|=0$. It follows that $\lambda\in F(A(\theta)+B(\theta))$. Hence we have $(-2\sin\theta,2\sin\theta)\subset F(A(\theta)+B(\theta))$.
We prove that $F(A(\theta)+B(\theta))\subset[-2\sin\theta,2\sin\theta]$. In Theorem 4 in [15], it is shown that $\sigma(A(\theta)+B(\theta))\subset[-2\sin\theta,2\sin\theta]\cup\{2\}$. Since $F(A(\theta)+B(\theta))\subset(0,2)$, we have $F(A(\theta)+B(\theta))\subset(0,2\sin\theta]$. ∎
Proof of (ii) of Example 5.4.
From statements (vi)-(x) of
Lemma A.1, we have
@@
Proof of (iii) of Example 5.4.
This follows directly from statement (i) and the fact that $F(\tau(g_{1})+\tau(g_{2}))$ is invariant under any unitary intertwining operator for any unitary representation $(\tau,V)$ of $D_{\infty}$. ∎
Proof of (iv) of Example 5.4.